Check for answers here first Hall A wiki Online HowTo, and return to this page if you can't find them.
First, do a visual survey of the Hall to make sure nothing is in the spectrometer's path. Also, make sure you are not exceeding the minimum (typically 14 degrees) or maximum (experiment dependent) spectrometer angle.
Second, on hacsbc2, click the red "tool box" icon on the linux taskbar, as above. Choose bogies_SetSpec so that you can determine the angle and vernier setting for the spectrometer. Enlarge See the window displayed to the right. Enter the spectrometer (L or R), and the angle, and you will get two options for the floor mark and the vernier. Generally choose the vernier closer to zero. Center the cameras on the desire vernier using the Move+/Move- buttons on the Hall A General Tools screen. The TV monitors for these cameras are on the middle shelf, in rack CH01A05.
Third, choose bogies_Left (or bogies_Right) in the tool box to bring up the bogies control screen. Click PSM enable and wait a few seconds for PSM OK to read YES. Click DM enable and wait a few seconds for DM OK to read YES.
Make sure the velocity is set to 0 and the direction is CW or CCW as desired. Click on Brake Release and wait for Brakes OK to read YES.
Click on ClampRelease, set the velocity to 700. Once you see the spectrometer start to move in the floor angle camera - you cannot see the spectrometer move in the Hall overview camera, as it only moves a few degrees per minute at maximum speed. For the left arm, to move to a larger angle, the direction should be CCW, while for the right arm CW moves the spectrometer to larger angle. The direction of the spectrometer is reversed by using a negative rpm. Watch the spectrometer motion on the cameras. When you are getting close to the desired angle, slow down to about 300 rpm. To stop, click on the Clamp Release button and the Brake button. Disable DM and PSM, and disconnect to close the GUI. Read off the floor angle mark and vernier, and input the values into the appropriate fields in the Alignment section of the Hall A General Tools GUI.
From the HAC window, there is a P0 SET field for each spectrometer in which you can enter the desired momentum setting. To change to a lower momentum from a higher previous setting, you just need to enter the new momentum value. To raise the momentum setting, you need to cycle the Q2 and Q3 magnets first and then enter the desired momentum setting. See the instructions for the cycling procedure on the instruction sheet to the right of the white board in the counting house.
New Momentum Control Instructions with information on:
* Momentum Regulation * Momentum Tolerance Alarms * Q2 & Q3 Degaussing * Troubleshooting
These instructions are taken from John Lerose's document. Due to the fact that there is a short in one of the coils of the right dipole there are special considerations that must be taken into account when setting it. The short imposes two limits on operational use of the magnet. The ramp rate is limited (~0.1 Amps sec-1) in order to keep the current induced in the short low enough to ensure safe operation of the magnet. This alone makes changing the field a slow process. Once current is induced in the short, then, left to itself, that current dies off exponentially, which affects the magnetic field, with a time constant close to 20 minutes. If one ignores this 2nd factor it can take a very long time to get a stable field in the magnet.
There is nothing, short of taking the magnet apart and fixing it, that can be done about #1 above. However, with a little patience and forethought, #2 can be dealt with fairly easily. The trick is to use the inductive coupling between the shorted and not shorted coil to bring the current in the short to zero. This is done by always overshooting the desired current change and then going back. Use Figure 1 to determine the amount of overshoot you need for a given change in current (Δi) or you can use the spreadsheet calculator.
For example, if the right dipole is set for 0.3616 GeV/c the output current is about 126.0 Amps (with a set current of about 123.8) and the desired field is -0.13268 T. To change to +10% higher, p = 0.3978 GeV/c, should require about 10% more current or 12.6 Amps more. Instead of just typing in iset = 136.4 (123.8+12.6), do the following:
Re the overshoot: Figure 1 is based on a model of the shorted coil with inductive coupling to the not shorted coil. As we get more experience we may refine that. In general for bigger changes more overshoot is needed. For changes less than 3 Amps just make the change and then make small adjustments. Also remember when going down the "overshoot" really is an undershoot. Always go past and then come back.
A final note: This assumes that the current in the short was zero when you started.
The reset for Q2 is simple, on the front of the power supply there are red lights on when the power supply is tripped. At the bottom of the board that has the red light there is a blue reset button. There may be more than one light on so there may be more than one blue button that needs to be pushed. Push blue buttons until all of the red lights are out. Once all of the lights are out, lower the large lever on the lower right side of the power supply and lift it back into position. this will reset the Q2 and Q3 power supplies locally. For Q1, it's simpler : there is a large sign in the back of the Q1 equipment rack on the power supply balcony that says the blue reset button is here. you press the button and the magnet resets.
HALOG entry 186341 by J. Segal: Sometimes the NMR meters need to be power cycled. Most likely when the meters do not respond to GPIB commands. To check if the meter is responding one can try to change the units on the NMR GUI. It's the GUI you get to from the "FIELDS" button and selecting "D1 NMR". Try changing units from "MHz" to "Tesla". If the units on the display change, the meter is responsive and does not need to be power cycled. If the units do not change it would be good to power cycle the meter.
(Username and password for IOC reboot are available in the counting house).
Halog entry 188657 by E. Brash: The IOC reboot is generally straightforward, but we had some problems getting the NMR to lock. I had some success with just setting the DAC manually to an appropriate value, and then turning the auto search on once it got close. I found that the following worked pretty well:
Right arm -> DAC=2400 Probe=B Left arm -> DAC=1319 Probe=B
Settings for NMR to lock:
Units : Tesla
Mode : Auto
Field : polarity of particule ( negative for electrons )
Search : Off
Go to the hall A main menu screen.
Click "IOCS". Note that the left HRS magnet IOC's (iocha14) are accessible at http://hareboot10 and the right HRS magnet IOC's(iocha16) are accessible at http://hareboot25 (Username and password for IOC reboot are available in the counting house). Close the "Hall A IOC Systems" window.
After the IOC reboot, please make sure that the spectrometer-angle information is reentered into the Hall A General Tools GUI.
VDC status can be checked on the HAC Main Control Window. High Voltages and Gas Flows are constantly being monitored in the VDC HV and Gas Flow sections.
From the Hall A Main Menu on hacsbc2, click on LeCroy HV. Choose the appropriate spectrometer, and then the appropriate detector. Usually the voltages are already plateaued, and you should only need to turn HV on. All voltage numbers are in Volts.
Bring up the HAC control window per the above instructions. In the "Miscellaneous" section, click on the "Crate Resets" button then select "Beamline Crates Resets". This will bring up the "Hall A Slow Controls" window. In the bottom left and right of this window are the reset buttons for the spectrometer subsystems. Click on the relevant "VDC high voltage" reset button. You then need to go back to the main HAC GUI and reset the tripped VDC to 4 kV by entering 4 in the VDC HV window. If the VDC trips again, reset the supply then try to ramp up the voltage slowly. If there are repeated trips, call an expert.
The main HV is provided by a LECROY 1440 crate for 253 channels of detectors. It is controlled via the serial interface which is accessible on hatsv4 port 6.
telnet hatsv4 2006
If you get a connection refused message, it means it is already connected. Login as root on the portserver
This will free port 6 of the portserver
A few useful commands :
M01 to select the crate, only needed at power up
R E A display all HV
ON to turn HV ON
OFF to turn HV OFF
ST to get status of crate
R E C[channel] read HV from the channel
R E A to display all values of HV, note there is a space between the letters
W [HV value] C[channel] Write the HV value
The rest of the HV is on the left arm HV crate in the CSR slots and can be controlled using the usual Hall A HV GUI
For easier communication with the HV crate, we have a small telnet server using Python which forward commands
to the crate.All software is in ../HV after a goonlana. ( /adaqfs/home/adaq/e05110/HV/c )
If the server is not running, launch it on logged adaql2 by typing in scripts/1440/HV/ :
The server is tcp, it will forward any command it receives to the crate.
A few simple C program are locate in the C directory, they allow to write or read a value and can serve as a base for more advanced program.
gethv [channel] display the HV of channel
chhv [channel] [HV value] write HV value in channel
loadhv [textfile] load values in the text file in the crate, format is channel number and requested value
comhv [1440 command] sends the command to the crate
channels 4 8 19 are bad and not used
channel for PMT 66 (HV 69)and 150 ( HV 152 ) are off since the PMT are bad
This assumes that runcontrol is running. If not, see the section Cold Start below.
* 2-daq mode: BigBite: The BigBite runs on adaql2 on the a-onl account. LeftHRS: The Left HRS runs on adaql1 on the adev account. * pedrun: Used to obtain pedestal files used by fastbus for pedestal suppression.
Generally adaql1 and adaql2 should be reserved for CODA / data acquisition. Replay is particularly resource intensive and will interfere with DAQ, but more generally nothing should be done on these machines except CODA and an ssh to bring up a halog window.
More details about the Hall-A DAQ and trigger can be found in Bob Michaels' guide_2daq and guide2_2daq .
If you've been running the normal spectrometer DAQ configuration, you'll need to press "Abort", then "RunTypes" and select the PEDRUN configuration. Download, StartRun. Run for about 5000 events, then EndRun. You may check the pedestal files in ~/ped (pedN.dat corresponding to ROC #N). See README there for more details. After pedestal run, change back to the running configuration.
NOTE: DON'T try to use PEDRUN for anything other than pedestal determination -- its confusing since the prescale factors are in a different file, etc. If you want a run with pedestal suppression turned off, see the README.
See the Bob Michael's guide2_2daq (section III, Quick Resets) for more details.
See the Bob Michael's guide2_2daq (section III, Cold Start) for more details.
xscaler is the GUI that displays scaler data online. Normally xscaler is running on the hapc5 screen which is above the adaql2 console (in rack CH01A09).
If not running:
The displayed scalers can be changed by clicking on the tabs at the top of the GUI. There are instructions at the bottom of the GUI explaining how to plot scaler channels versus time and to switch the display from rates to counts.
Usually if there is a problem, it is noticed because no data is coming in, or because the system fails to stop or start runs. Check that the ROCs (Read Out Controllers) are up and if not reboot them. For more information see Bob Michaels DAQ guide_2daq. If you are not a coda expert, it might be fastest to simply exit the coda runControl window, and then restart coda as above.
To change the trigger:
A self-explanatory graphical user interface pops up and you should do what it says and press "Download" and WAIT for it to finish and do not press Ctrl-C. The user should look for suspicious error messages in the window from which trigsetup was launched, e.g. to check if connection to the crate is ok.
If it is not already running, type datamon in a terminal on adaql2 in the a-onl account for BigBite, and on adaql1 in the adev account for the LHRS.
As adev on adaql1 or a-onl on adaql2 type :
This will open 3 xterms for either the BigBite or left DAQ, respectively. In the name of these xterms is the portserver port where you can connect via RS232 (portserver instructions are at hallaweb.jlab.org/equipment/daq/portserver.html).
In the "components" workspace, login to all the frontend computers by typing "telnet hallasfi1" where the name of the computer like hallasfi1 is in the name of the xterm window. An example name may be an xterm with title "ROC2--hallasfi2--hatsv3-port-8" meaning ROC2 is IP address hallasfi2 and on portserver hatsv3 at port 8.
For example to log on hatsv3 port 8 type :
telnet hatsv3 2008
Use spot++ as described below. Also, the raster pattern should be displayed on the Tektronix oscilloscope mounted in rack CH01B05, in the middle room of the counting house. This can give you a rough idea that the raster is on, but spot++ is needed to check actual dimensions. Also see this page for more.
How to start spot++
* Start a CODA run on L-arm or R-arm DAQ * Log onto "adaql1" as user "adaq" * Type "spot_L" ("spot_R") for L-arm (R-arm)
A few notes:
MORE DETAILS ON USAGE:
* Type "spot_L" or "spot_R" with no arguments --> to analyze the most recent run. * Type "spot_L 1062" --> to analyze run 1062 (if it's still on adaq disks). * Type "spot help" --> to get help (note, there are other command line options, which "help" will explain).
Please see http://hallaweb.jlab.org/equipment/daq/spot.html for more details.
Please follow the directions at the Harp Scan Page (Need better instructions!) [Harp Scan page]
The machine which runs the HAC GUI is hacsbc2. The monitor is located in the Hall A Counting House in rack CH01A02, just below the NMR dipole oscilloscopes. Click here to see the screen. At the bottom of the screen is the linux taskbar, in the middle of the taskbar is the "toolbox" icon. Pressing on that icon will open a menu with the following choices : vdcs_off, T2.stp, StripTool, Snapshot, etc; as shown on the picture aside. Click the entry called "Menu_HallA", or "hlamain". You can also issue the command "hlamain" from the terminal.
A sub-screen entitled "Hall A Main Menu" will appear. One of the entries is labeled "Tools". Clicking on the "Tools" button opens the "Hall A General Tools" screen, which is often referred to as the HAC window.
From the "toolbox" icon in the taskbar, press the "Alarm Handler" label. Alarms are indicated by a beep and the alarm handler button flashing. The color red indicates values far out of range, while the color white indicates a communication failure. Clicking on the alarm handler button will bring up an alarm handler list. A detailed list is given in the left panel, and clicking on items here brings up a more detailed list in the right panel. The alarm can be acknowledged and silenced by clicking on the button beside the alarm indicator.
Pressing the "toolbox" icon and launch "StripTool", which enables you to plot any EPICS variable(s) versus time. To find out the name of a particular EPICS variable, click on it with the middle mouse button on the MEDM GUI.
See http://hallaweb.jlab.org/equipment/daq/abu.html for more information.
If no halog guis are up and running, log in as user "adaq" on any "adaq.." machine. Type halog a yellow and blue and gray window will appear Enter text in the large message area, your name in the User area, and a title in the Keywords area. Remember that experiments tend to suffer from too little documentation, not too much. Also, halog entries allow people to remotely check on the experiment progress without bothering you with phone calls. When complete, after pressing the "Make entry" button, the text in the message center to the lower right corner of the window should read "Entry Complete", otherwise something is wrong. It often takes several minutes for entries to show up in halog.
To enter graphics: Click "Grab Screen" in the bottom button row of the halog window. The halog window will disappear, and the window the cursor is in will become the active window, its outline flashing. By moving the cursor you can select any window to be selected, by clicking on the left mouse button. You also can select an arbitrary part of the screen by holding down the left mouse button and dragging the cursor along. All the time a rectangle will flash that indicates the selected area. Releasing the mouse button will then cause this screen area to be included into the entry. It will show up as a thumbnail picture in the halog window, where you can discard it or accept it, and then make a second screen grab. The images are in GIF format.
Online Run analysis:
See https://hallaweb.jlab.org/wiki/index.php/Transversity/Online_Replay for more information.
Please see the data structure page: http://hallaweb.jlab.org/equipment/daq/dstruct.html
The target temperature and pressure is displayed on the target GUIs running on cryotarg as well as on the camera monitors in the rack (CH01A06) left of the target computer. Look at the camera monitor on the top right of rack CH01A06. It displays the temperatures for the three loops on three white temperature controllers, with loop 1 at the top and loop 3 on the bottom. Ask the TO for help if you can't find the temperature or it doesn't make sense. The adjacent camera displays the JT valve controller readings and tachometers.